If a1 a2 a3.... an are in ap with a1= 0 then (a3/a2 + a4/a3 .......an/an-1 ) - a2(1/a2 +1/a3 .... 1/an-2

Grade:10

1 Answers

Deevena
3 years ago
a/1-b=b/1-c=c/1-c=k
a=k-kc
C=k-k^2+k^2(k-k c) 
a=k-k^2+k^3-k^3c
a=k-k^2+k^3/1+k^3=b=c
but a₦b₦c
I. e. K^3=-1=>k=-w, -w^2